# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated surprising performance across various natural language processing tasks. Recently, medical LLMs enhanced with domain-specific knowledge have exhibited excellent capabilities in medical consultation and diagnosis. These models can smoothly simulate doctor-patient dialogues and provide professional medical advice. Most medical LLMs are developed through continued training of open-source general LLMs, which require significantly fewer computational resources than training LLMs from scratch. Additionally, this approach offers better protection of patient privacy compared to API-based solutions. This survey systematically explores how to train medical LLMs based on general LLMs. It covers: (a) how to acquire training corpus and construct customized medical training sets, (b) how to choose a appropriate training paradigm, (c) how to choose a suitable evaluation benchmark, and (d) existing challenges and promising future research directions are discussed. This survey can provide guidance for the development of LLMs focused on various medical applications, such as medical education, diagnostic planning, and clinical assistants.
Index Terms:
Large Language Model, Domain, Medicine, Healthcare.
I Introduction
@@
Recently, large language models (LLMs), such as ChatGPT [1], have demonstrated exceptional capabilities in handling natural language processing (NLP) tasks. Their performance improves significantly as the scaling of the model increases, while the size of the dataset keeps expanding. Currently, LLMs are widely applied in intelligent education, legal consultation, code generation, healthcare, and finance, showcasing their powerful NLP capabilities and broad applicability. In intelligent education [2, 3], LLMs find application in intelligent tutoring systems and personalized learning planning, providing personalized advice by analyzing students’ learning behaviors and feedback. In legal consulting [4, 5], LLMs facilitate legal document analysis, contract review and legal question answering, improving the efficiency and accuracy of legal services. In code generation [6, 7], LLMs can automatically generate code snippets, perform code completions and fixes, augmenting development efficiency and code quality. In heathcare [8, 9], LLMs are used for medical literature analysis and doctor-patient dialogues, assisting medical decisions and improving patient care. In finance [10, 11], LLMs are employed for risk assessment, financial analysis and customer service, enhancing the intelligence and personalization of financial services.
@@
The success of these above domain LLMs relies on inheriting the strong knowledge of general LLMs and performing further training injected with domain-specific knowledge. The general LLM has acquired extensive knowledge, linguistic styles, and generalization abilities across a diverse corpus, enabling it to quickly learn new tasks. Therefore, the strategy of training domain LLMs based on general LLMs is more economical and effective than training domain LLMs from scratch [12]. There are three optional training phases for training the medical LLM from the general LLM [13, 14, 15]: Continued Pretraining (CP), Instruction Fine-tuning (IFT) and Human Alignment (HA). CP involves training on unstructured domain data to help LLMs learn domain-specific knowledge, terminology, and language style. IFT focuses on training with domain instruction data, enabling LLMs to master domain-specific dialogue and instruction following. HA trains LLMs on human preference data, equipping it with domain-specific qualities such as harmlessness and helpfulness.
@@
Health constitutes the cornerstone of human survival and development, and advancements in medical technology and services significantly impact quality of life and longevity [16, 17]. Given the aging population and rising prevalence of chronic diseases, the medical domain confronts disparities in resource allocation, necessitating more efficient solutions [18]. Medical LLMs can address this challenge by integrating rich medical data and clinical cases to help physicians quickly and accurately make diagnoses and formulate treatment plans [19]. Furthermore, it can help medical institutions optimize resource allocation and improve the efficiency and quality of medical services. Existing medical LLMs focusing on different medical disciplines, such as internal medicine, respiratory medicine, and gastroenterology, can provide more specialized answers compared to general LLMs. Additionally, in medical consultation tasks, medical LLMs excel in interactive diagnosis, proactively prompting and guiding patients to provide missing information for ambiguous descriptions, distinguishing them from general LLMs.
@@
Figure 1: Training Pipeline from General LLMs to Medical LLMs. Firstly, the medical corpus is collected and processed to form a standardized training set. Next, an appropriate training paradigm is selected to train General LLMs to become Medical LLMs with medical knowledge. The training paradigms consist of three optional training phases: Continued Pretraining (CP), Instruction Fine-tuning (IFT), and Human Alignment (HA). Finally, Medical LLMs are evaluated from both machine and human perspectives.
@@
Given the significance of the healthcare domain, the advantages of medical LLMs over general LLMs, and the cost-effectiveness and efficiency of training domain LLMs from general LLMs, this paper systematically surveys medical LLMs derived from further training of general LLMs from the perspectives of dataset, methodology and evaluation. The main content is depicted in Fig. 1. Firstly, we focus on the construction of training datasets for medical LLMs, as the quality of these datasets directly impacts the models’ medical capabilities. Thus, we analyze the training datasets for existing medical LLMs from two perspectives: the source of the training corpus and data processing methods. Training corpus sources can assist developers in collecting the expected data, while data processing methods facilitate the cleansing of data to form a standardized training set. Furthermore, considering the differing capabilities injected to medical LLMs by the three training stages of CP, IFT, and HA, as well as the variations in required computational resources and data scale, we have summarized the combinations of training stages adopted by existing medical LLMs and classified them into four paradigms: IFT, CP-IFT, IFT-HA, and CP-IFT-HA. These paradigms can guide institutions in choosing the appropriate training methods based on their computational capabilities, data scale, and specific needs. Subsequently, given the professional nature of content generated by medical LLMs and the potential harm that erroneous content can cause to patients, we systematically review existing evaluation benchmarks and protocols from two perspectives: machine evaluation and human-centric evaluation. In particular, we standardize the dimensions of the human-centric evaluation. Finally, we analyze the shortcomings of existing medical LLMs and propose future research directions with the potential to address these issues.
@@
Here, we compare several recent similar studies. The study [20] systematically summarizes the training data, methods, optimization strategies, and evaluation techniques for LLMs in healthcare. However, our work retrieved updated literature on medical LLMs and summarized the training data and methods from a more nuanced perspective. Specifically, we analyzed the data sources and processing techniques in detail for training data. For training methods, we identified four training paradigms and examined their respective advantages and disadvantages. Another study similar to ours is [21], which systematically overviews the methods, datasets, and applications details of language foundation models, vision foundation models, bioinformatics foundation models, and multimodal foundation models in healthcare. Our work specifically focuses on language foundation models, summarizing datasets, training methods, and evaluation benchmarks using new classifications and perspectives. Finally, some studies [19, 22] have concentrated on the applications and challenges of LLMs in medicine. In contrast, our research specifically emphasizes methods and technologies.
The contributions of this review are summarized below:
@@
    We systematically review medical LLMs emerged since ChatGPT. Detailed guidelines and tutorials are provided for healthcare organizations in training customized medical LLMs in terms of dataset construction, training technologies and evaluation benchmarks.
@@
    Data Acquisition and Processing (Section II). Summarizing and classifying existing training corpus sources and data processing methods for medical LLMs offers robust recommendations and guidance for constructing customized medical datasets.
@@
    Training Paradigms (Section III). Systematically summarizes the existing training methods for medical LLMs and proposes four distinct training paradigms. This facilitates medical organizations in selecting appropriate training methods based on their available computational resources.
@@
    Medical LLMs Evaluations (Section IV). Summarizes the existing benchmarks for evaluating medical LLMs and categorizes them into machine and human-centric evaluations. Additionally, the human-centric evaluation dimensions of medical LLMs are normalized. This provides a comprehensive assessment perspective and methodology for healthcare organizations.
@@
    Challenges and Future Directions (Section V). Future research directions are proposed to address the shortcomings and research gaps of the existing medical LLMs. This provides important guidance and insights to further exploration in Medical LLMs for medical organizations and researchers.
II Data Acquisition and Processing
Datasets are crucial for transferring knowledge of LLMs from general domains to specific ones. The quality of a dataset significantly impacts model performance. A high-quality dataset provides accurate supervised signals, enabling the model to effectively learn domain-specific rules and patterns. This not only improves the model’s alignment with the target task but also break the scaling law to some extent [23]. Medical LLMs should be doctor-like, patient-friendly, and professional [8]. To ensure that these models are able to learn the language characteristics and terminology of the medical domain, training datasets should meet the above guidelines. To construct a medical domain dataset that meets the requirements of a specific task, it is essential to identify appropriate training corpus sources. These sources may vary in their medical discipline focus, level of specialization, and data quality. Therefore it is crucial to select a corpus source that meets the needs. After identifying a suitable corpus source, the corpus cannot be directly used for model training. They may contain inappropriate content, such as violent, pornographic, or discriminatory material, or have data format irregularities. Therefore, data processing is required to ensure the corpus meets the necessary training requirements. The training of medical LLMs currently requires three types of data: unstructured data for continued pretraining, question-answer pairs (both single and multiple turns) for instruction fine-tuning, and human preference data for human alignment. Here, we systematically summarize the training corpus sources and data processing methods for investigated medical LLMs in Tab. I, providing a reference for constructing custom medical training sets. In addition, we categorized the corpus sources and data processing methods, as shown in Fig. 2.
@@ II Data Acquisition and Processing
TABLE I: Corpus Sources, Processing Methods, Training Set Format and Scale for Medical LLMs. The abbreviations here are Pub. Data: Existing Public Datasets, Pub. Corpus: Public Medical Corpus, Org. Corpus: Professional Medical Organization Corpus , Syn.:Synthetic Data, clean.: Data Cleaning, Format.: Data Formatting & Construction, Aug.: Data Augmentation, Trans.:Translation, QA: Question-Answer datasets (single turn or multiple turns or multiple choice), UD: Unstructured Data, HPD: Human Preference Data, KG: Knowledge Graph, NLP: NLP Task, MB: Medical Website, Encyc.: Encyclopedia, TB: Textbook, Guid.: Guideline, PI: Package Insert, MAC: Medical Academic Literature, CN: Clinical Note, EHR: Electronic Health Record, MP: Medical Prescription, DPC: Doctor-Patient Conversation
@@
Models Training Corpus Sources Data Processing Standardized Training Sets
Pub. Data. Pub. Corpus Org. Corpus Syn. Clean. Format. Aug. Trans. Format (Scale)
Med-PaLM [24] QA MB InsD (65 items)
ChatDoctor [25] MB $\checkmark$ InsD (100K items)
DoctorGLM [26] $\checkmark$ InsD (4,487K items)
BenTsao [27] KG $\checkmark$ InsD (8K items)
ChatGLM-Med [28] KG $\checkmark$ InsD (7.6K items)
MedAlpaca [29] QA, NLP MB $\checkmark$ $\checkmark$ InsD (248K items)
PMC-LLaMA [30] UD, QA, KG MAL, TB $\checkmark$ $\checkmark$ $\checkmark$ UD (79B tokens),
InsD (202M tokens)
HuatuoGPT [8] QA DPC $\checkmark$ $\checkmark$ $\checkmark$ InsD (226K items)
ChatMed-Consult [31] MB $\checkmark$ $\checkmark$ InsD (110K items)
Med-PaLM 2 [32] QA InsD (193K items)
Clinical Camel [33] QA MAL $\checkmark$ $\checkmark$ InsD (174K items)
ShenNong-TCM [34] KG $\checkmark$ InsD (113K items)
MedicalGPT [35] QA, HPD UD (361K items),
InsD (2.07M items),
HPD (3.8K items)
ClinicalGPT [36] QA, KG EHR InsD (423K items),
HPD (10K items)
DISC-MedLLM [37] QA, KG $\checkmark$ $\checkmark$ InsD (514K items)
Zhongjing [38] QA, KG,
NLP
TB EHR, CN,
@@
$\checkmark$ $\checkmark$ UD (1,086 MB),
InsD (470K items),
HPD (20K items)
BianQue [39] DPC $\checkmark$ $\checkmark$ InsD (2,437K items)
Alpacare [40] $\checkmark$ $\checkmark$ InsD (52K items)
@@
$\checkmark$ UD (2B tokens),
InsD (11M items),
HPD (6K items)
Taiyi [42] QA, NLP $\checkmark$ $\checkmark$ InsD (1,114K items)
ChiMed-GPT [43] UD, QA, HPD $\checkmark$ $\checkmark$ UD (369K items),
InsD (1,269K items),
HPD (4K items)
MediTron [44] UD, QA MAL CN $\checkmark$ $\checkmark$ UD (48 tokens),
InsD (369K items)
HuatuoGPT-II [45] UD Encyc., TB, MAL, MW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ InsD (5,394K items)
AntGLM-Med [46] UD, QA, KG TB, MAL DPC $\checkmark$ $\checkmark$ UD (15.39B tokens),
InsD (632K items)
GPT-doctor [47] MB, PI $\checkmark$ InsD (1,939K items)
BioMistral [48] UD, QA UD (3B tokens),
InsD (405K items)
MMedLM [49] UD MW, TB $\checkmark$ UD (25.5B tokens)
InMD-X [9] MAL $\checkmark$ $\checkmark$ UD (150M tokens),
InsD (1,701K items)
@@
MAL EHR, CN $\checkmark$ UD (129B tokens),
InsD (214K items)
JMLR [51] QA TB, Guid. EHR UD (-), InsD (-)
BiMediX [52] QA $\checkmark$ $\checkmark$ InsD (1,311K items)
OncoGPT [53] MB DPC $\checkmark$ InsD (332K items)
Apollo [54] UD, QA TB, MAL, MW, Encyc., Guid. $\checkmark$ $\checkmark$ UD (2,054M tokens),
InsD (481M items)
Qibo [55] QA, NLP TB, Encyc. MP $\checkmark$ UD (1,277MB), InsD(-)
II-A Training Corpus Sources
We categorized the corpus sources into four groups: Existing Public Datasets, Public Medical Corpus, Professional Medical Organization Corpus, and Synthetic Data. The strengths and weaknesses of each category, as well as their suitability for different training phases of medical LLMs, are detailed below. Additionally, guidance is provided for customizing the collection of training data.
II-A1 Existing Public Datasets
It is recommended to use existing public datasets for constructing customized training sets for medical LLMs. These datasets are well-designed, high-quality, and widely recognized in existing research. The public datasets are further categorized into five types: unstructured data, question-answer pairs, human preference data, knowledge graphs, and regular NLP task data. Related work [20, 21] have systematically summarized datasets for LLMs in the medical domain. Here, we focus on the characteristics of each type of data and considerations for using these public datasets.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated surprising performance across various natural language processing tasks. Recently, medical LLMs enhanced with domain-specific knowledge have exhibited excellent capabilities in medical consultation and diagnosis. These models can smoothly simulate doctor-patient dialogues and provide professional medical advice. Most medical LLMs are developed through continued training of open-source general LLMs, which require significantly fewer computational resources than training LLMs from scratch. Additionally, this approach offers better patient privacy protection than API-based solutions. Given the above advantages, this survey systematically summarizes how to train medical LLMs based on open-source general LLMs from a more fine-grained perspective. It covers (a) how to acquire training corpus and construct customized medical training sets, (b) how to choose an appropriate training paradigm, (c) how to choose a suitable evaluation benchmark, and (d) existing challenges and promising research directions are discussed. This survey can provide guidance for the development of LLMs focused on various medical applications, such as medical education, diagnostic planning, and clinical assistants. Related resources and supplemental information can be found on the GitHub repository  [^1] https://github.com/jqwangai/Medical-LLM.
keywords:
[…]
[fn1]https://github.com/jqwangai/Medical-LLM
1 Introduction
@@
Recently, large language models (LLMs), such as ChatGPT [2], have demonstrated exceptional capabilities in handling natural language processing (NLP) tasks. Their performance improves significantly as the model scaling increases while the size of the dataset keeps expanding. Currently, LLMs are widely applied in intelligent education, legal consultation, code generation, healthcare, and finance, showcasing their powerful NLP capabilities and broad applicability. In intelligent education [32, 154], LLMs find application in intelligent tutoring systems and personalized learning planning, providing personalized advice by analyzing students’ learning behaviors and feedback. In legal consulting [31, 28], LLMs facilitate legal document analysis, contract review, and legal question answering, improving the efficiency and accuracy of legal services. In code generation [196, 35], LLMs can automatically generate code snippets and perform code completions and fixes, augmenting development efficiency and code quality. In healthcare [188, 53], LLMs are used for medical literature analysis and doctor-patient dialogues, assisting medical decisions and improving patient care. In finance [92, 82], LLMs are employed for risk assessment, financial analysis, and customer service, enhancing the intelligence and personalization of financial services.
@@
The success of these above domain LLMs relies on inheriting the strong knowledge of general LLMs and performing further training injected with domain-specific knowledge. The general LLM has acquired extensive knowledge, linguistic styles, and generalization abilities across a diverse corpus, enabling it to learn new tasks quickly. Therefore, the strategy of training domain LLMs based on general LLMs is more economical and effective than training domain LLMs from scratch [97]. There are three optional training stages for training the domain LLM from the general LLM [199, 114, 52]: Continued Pretraining (CP), Instruction Fine-tuning (IFT), and Human Alignment (HA). CP involves training on unstructured domain data to help LLMs learn domain-specific knowledge, terminology, and language style. IFT focuses on training with domain instruction data, enabling LLMs to master domain-specific dialogue and instruction following. HA trains LLMs on human preference data, equipping them with domain-specific qualities such as harmlessness and helpfulness.
@@
Health constitutes the cornerstone of human survival, and advancements in medical technology significantly impact the quality of life and longevity [194, 124, 197]. Given the aging population and rising prevalence of chronic diseases, the medical domain confronts disparities in resource allocation, necessitating more efficient solutions [135, 21]. Medical LLMs can address this challenge by integrating rich medical data and clinical cases to help physicians quickly and accurately make diagnoses and formulate treatment plans [141]. Furthermore, it can help medical institutions optimize resource allocation and improve the efficiency and quality of medical services. Existing medical LLMs focusing on different medical disciplines, such as internal medicine, respiratory medicine, and gastroenterology, can provide more specialized answers than general LLMs. Additionally, in medical consultation tasks, medical LLMs excel in interactive diagnosis, proactively prompting and guiding patients to provide missing information for ambiguous descriptions, distinguishing them from general LLMs.
@@
Figure 1: Training Pipeline from General LLMs to Medical LLMs. Firstly, the medical corpus is collected and processed to form a standardized training set. Next, an appropriate training paradigm is selected to train General LLMs to become Medical LLMs with medical knowledge. The training paradigms consist of three optional training stages: Continued Pretraining (CP), Instruction Fine-tuning (IFT), and Human Alignment (HA). Finally, Medical LLMs are evaluated from both machine and human perspectives.
@@
Given the significance of the healthcare domain, the advantages of medical LLMs over general LLMs, and the cost-effectiveness and efficiency of training domain LLMs from general LLMs, this paper systematically surveys medical LLMs derived from further training of general LLMs from the perspectives of dataset, methodology, and evaluation. The main content is depicted in Fig. 1. Firstly, we focus on the construction of training datasets for medical LLMs, as the quality of these datasets directly impacts the models’ medical capabilities. Thus, we analyze the training datasets for existing medical LLMs from two perspectives: the source of the training corpus and data processing methods. Training corpus sources can assist developers in collecting the expected data, while data processing methods facilitate the cleansing of data to form a standardized training set. Furthermore, considering the differing capabilities injected into medical LLMs by the three training stages of CP, IFT, and HA, as well as the variations in required computational resources and data scale, we summarized the combinations of training stages adopted by existing medical LLMs and classified them into four paradigms: IFT, CP-IFT, IFT-HA, and CP-IFT-HA. These paradigms can guide institutions in choosing the appropriate training methods based on their computational capabilities, data scale, and specific needs. Subsequently, given the professional nature of content generated by medical LLMs and the potential harm that erroneous content can cause to patients, we systematically review existing evaluation benchmarks and protocols from two perspectives: machine evaluation and human-centric evaluation. In particular, we standardize the dimensions of the human-centric evaluation. Finally, we analyze the shortcomings of existing medical LLMs and propose future research directions with the potential to address these issues.
@@
Here, we compare several recent similar surveys. The survey [58] systematically summarizes the training data, methods, optimization strategies, and evaluation techniques for LLMs in healthcare. However, it lacks sufficient depth in discussing data sources and evaluation techniques, and overlooks the summary of data processing methods. Compared to it, our work retrieves the latest literature and analyzes the training data from the perspective of 16 corpus sources. Additionally, our work analyzes data processing methods and evaluation techniques at a more fine-grained level. Another study similar to ours is [59], which systematically overviews the methods, datasets, and application details of language foundation models, vision foundation models, bioinformatics foundation models, and multimodal foundation models in healthcare. However, it lacks an in-depth analysis of both the data, training strategies, and evaluation techniques for the language foundation models. Some other surveys [141, 16] focused on the applications and challenges of the LLMs in medicine, rather than on methods and technology.
To bridge these gaps, the survey makes the following contributions:
@@
    We systematically survey medical LLMs that emerged since ChatGPT. Detailed guidelines and tutorials are provided for healthcare researchers in training customized medical LLMs in terms of dataset construction, training paradigms, and evaluation technologies.
@@
    Data Acquisition and Processing (Section 2). We summarize and classify 16 training corpus sources and 4 groups of data processing methods for medical LLMs, providing robust recommendations and guidance for constructing customized medical datasets.
@@
    Training Paradigms (Section 3). We systematically summarize the training stages for existing medical LLMs and classify them into 4 training paradigms. This categorization facilitates medical researchers in selecting appropriate training paradigms based on their available computational resources.
@@
    Medical LLMs Evaluations (Section 4). We summarize the existing benchmark types for evaluating medical LLMs and categorize them into machine and human-centric evaluations. Additionally, the human-centric evaluation dimensions of medical LLMs are normalized. This result provides a comprehensive evaluation perspective and methodology for healthcare researchers.
@@
    Challenges and Future Directions (Section 5). We propose future research directions for addressing the shortcomings and research gaps of existing medical LLMs. These directions provide important guidance and insights for medical organizations and researchers to explore medical LLMs further.
2 Data Acquisition and Processing
@@
Data are crucial for transferring knowledge of LLMs from general purpose to domain applications. The quality of a dataset significantly impacts model performance. A high-quality dataset provides accurate supervised signals, enabling the model to effectively learn domain-specific rules and patterns. Moreover, this high-quality dataset not only improves the model’s alignment with the target task but also breaks the scaling law to some extent [50]. Medical LLMs should be doctor-like, patient-friendly, and professional [188]. To ensure that general LLMs are able to learn the language characteristics and terminology of the medical domain, training datasets should meet the above guidelines. To construct a medical domain dataset that meets the requirements of a specific task, it is essential to identify appropriate training corpus sources. These sources may vary in their medical discipline focus, level of specialization, and data quality. After identifying suitable corpus sources, the corpus cannot be directly used for model training. The data may contain inappropriate content, such as violent, pornographic, discriminatory material, and irregular data formats. Therefore, data processing is required to ensure the corpus meets the necessary training requirements.
The training of medical LLMs requires three types of data: unstructured data for continued pretraining, question-answer pairs (both single and multiple turns) for instruction fine-tuning, and human preference data for human alignment. Here, we systematically summarize the training corpus sources and data processing methods for investigated medical LLMs in Tab. LABEL:tab:dataset, providing a reference for constructing custom medical training sets. In addition, we categorized the corpus sources and data processing methods, as shown in Fig. 2.
2.1 Training Corpus Sources
We categorized the corpus sources into 4 main groups: Existing Public Datasets, Public Medical Corpus, Professional Medical Organization Corpus, and Synthetic Data. These groups were further subdivided into 16 categories, as illustrated in Fig. 2. The strengths and weaknesses of each category, as well as their suitability for different training stages of medical LLMs, are detailed below. Additionally, the guidance is provided for the collection of customizing training data. Finally, we counted the frequency of each corpus source according to the data in Tab. LABEL:tab:dataset, as shown in Fig. 3.
@@ 2.1 Training Corpus Sources
Table 1: Corpus Sources, Processing Methods, Training Set Format and Scale for Medical LLMs. The abbreviations here are Pub. Data: Existing Public Datasets, Pub. Corpus: Public Medical Corpus, Org. Corpus: Professional Medical Organization Corpus, Syn.:Synthetic Data, clean.: Data Cleaning, Format.: Data Formatting & Construction, Aug.: Data Augmentation, Trans.:Translation, QA: Question-Answer datasets (single turn or multiple turns or multiple choice), UD: Unstructured Data, InsD: Instruction Data, HPD: Human Preference Data, KG: Knowledge Graph, NLP: NLP Task, MB: Medical Website, Encyc.: Encyclopedia, TB: Textbook, Guid.: Guideline, PI: Package Insert, MAC: Medical Academic Literature, CN: Clinical Note, EHR: Electronic Health Record, MP: Medical Prescription, DPC: Doctor-Patient Conversation.
@@
Models Training Corpus Sources Data Processing Standardized Training Sets (Size)
Pub. Data. Pub. Corpus Org. Corpus Syn. Clean. Format. Aug. Trans. UD InsD HPD
Med-PaLM [130] QA MW 65 items
ChatDoctor [91] MW $\checkmark$ 100K items
DoctorGLM [172] $\checkmark$ 4,487K items
BenTsao [151] KG $\checkmark$ 8K items
ChatGLM-Med [55] KG $\checkmark$ 7.6K items
MedAlpaca [54] QA, NLP MW $\checkmark$ $\checkmark$ 248K items
PMC-LLaMA [167] UD, QA, KG MAL, TB $\checkmark$ $\checkmark$ $\checkmark$ 79B tokens 202M tokens
HuatuoGPT [188] QA DPC $\checkmark$ $\checkmark$ $\checkmark$ 226K items
ChatMed-Consult [164] MW $\checkmark$ $\checkmark$ 110K items
Med-PaLM 2 [131] QA 193K items
Clinical Camel [144] QA MAL $\checkmark$ $\checkmark$ 174K items
ShenNong-TCM [163] KG $\checkmark$ 113K items
MedicalGPT [174] UD, QA, HPD 361K items 2.07M items 3.8K items
ClinicalGPT [150] QA, KG EHR 423K items 10K items
DISC-MedLLM [11] QA, KG $\checkmark$ $\checkmark$ 514K items
Zhongjing [179] QA, KG, NLP TB EHR, CN,
@@
$\checkmark$ $\checkmark$ 1,086 MB 470K items 20K items
BianQue [23] DPC $\checkmark$ $\checkmark$ 2,437K items
Alpacare [195] $\checkmark$ $\checkmark$ 52K items
@@
$\checkmark$ 2B tokens 11M items 6K items
Taiyi [102] QA, NLP $\checkmark$ $\checkmark$ 1,114K items
ChiMed-GPT [142] UD, QA, HPD $\checkmark$ $\checkmark$ 369K items 1,269K items 4K items
MediTron [24] UD, QA MAL CN $\checkmark$ $\checkmark$ 48 tokens 369K items
HuatuoGPT-II [22] UD Encyc., TB, MAL, MW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 5,394K items
AntGLM-Med [89] UD, QA, KG TB, MAL DPC $\checkmark$ $\checkmark$ 15.39B tokens 632K items
GPT-doctor [155] MW, PI $\checkmark$ 1,939K items
EpilepsyLLM [200] QA MW MP, CN $\checkmark$ $\checkmark$ 52.2K items
BioMistral [79] UD, QA 3B tokens 405K items
MMedLM [121] UD MW, TB $\checkmark$ 25.5B tokens
InMD-X [53] MAL $\checkmark$ $\checkmark$ 150M tokens 1,701K items
@@
MAL EHR, CN $\checkmark$ 129B tokens 214K items
JMLR [153] QA TB, Guid. EHR - -
BiMediX [118] QA $\checkmark$ $\checkmark$ 1,311K items
OncoGPT [69] MW DPC $\checkmark$ 332K items
Apollo [156] UD, QA TB, MAL, MW, Encyc., Guid. $\checkmark$ $\checkmark$ 2,054M tokens 481M items
Qibo [189] QA, NLP TB, Encyc. MP $\checkmark$ 1,277MB -
Hippocrates [3] UD, QA $\checkmark$ 298M tokens 182.9M tokens 15,258 items
MING-MOE [94] QA, NLP $\checkmark$ 300K tokens
LingDan [67] UD PI, TB MP, CN, EHR $\checkmark$ $\checkmark$ 304M tokens 201.5K items
Aloe [51] QA Guid. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 872K items 12k items
PediatricsGPT [177] QA. KG TB, Encyc., Guil. DPC $\checkmark$ $\checkmark$ 975.8MB 332K items 15,556 items
Aqulia-Med [198] UD, QA $\checkmark$ $\checkmark$ $\checkmark$ 80B tokens 320K items 12,727 items
Figure 3: Percentage of frequency for each corpus source. The abbreviations are the same as in Tab. LABEL:tab:dataset.
2.1.1 Existing Public Datasets
It is recommended that existing public datasets be used to construct customized training sets for medical LLMs. These datasets are well-designed, high-quality, and widely recognized in existing research. The public datasets are further categorized into five types: unstructured data, question-answer pairs, human preference data, knowledge graphs, and regular NLP task data. Related work [58, 59] have systematically summarized datasets for LLMs in the medical domain. Here, we focus on the characteristics of each data type and considerations for using these public datasets.